Mathematics 242 – Principles of Analysis Definitions Quiz 7 – April 8, 2011

Your Name:

A) (10) Let f be defined on an open interval containing x = c. What does it mean to say f is differentiable at c?

B) (10) What does it mean for a set S ⊂ R to be countably infinite?